OpenStudy (anonymous):

solve for x-3y=5 and -4x+6y=-15

OpenStudy (anonymous):

What do you think the first step is to do? You can substitute/eliminate.

OpenStudy (anonymous):

I can solve for one of the equations for x-3y=15

OpenStudy (anonymous):

So you want to do substitution? Where you have what one variable equals and you plug it into the other equation? Such as x=some_numbers In this case x-3y=15 You add the 3y to the other side so you have x=16+3y Do you know what the next step is?

OpenStudy (anonymous):

I would plug it in to the other equarion would it be -4{16+3y)+6y=-15

OpenStudy (anonymous):

Correct! Now what does that distribute out to equal?

OpenStudy (anonymous):

Y= -49/6?

OpenStudy (anonymous):

Oh, I'm sorry I made a typo! In this case x-3y=15 You add the 3y to the other side so you have* x=16+3y* <- that should equal x=5+3y

OpenStudy (anonymous):

so you get -4(5+3y)+6y=-15

OpenStudy (anonymous):

I tried figuring it out would the answer be (38/5,-5/6)?

OpenStudy (anonymous):

-5/6 is correct but the other one isn't, where do you plug y=-5/6 back into, to solve for x?

OpenStudy (anonymous):

Would it be 5/2 i would plug in -5/6 to x-3y=5

OpenStudy (anonymous):

Yup, I'm pretty sure that is correct!